Bethany invested $2500, part at 8% per year and the rest at 12% interest per year. In one year, the parts earned equal amounts of interest. determine how much was invested at each rate.
Answer by stanbon(75887)

Bethany invested $2500, part at 8% per year and the rest at 12% interest per year. In one year, the parts earned equal amounts of interest. determine how much was invested at each rate.
-------------------
Let amount invested at 8% be "x" ; interest on this is 0.08x dollars
----------------
Amt left to invest is "2500-x"; interest on this is 0.12(2500-x) dollars
-----------------
EQUATION:
1st interest = 2nd interest
0.08x = 0.12(2500-x)
8x = 12(2500-x)
8x = 30000-12x
20x =30000
x = 1500 dollars (amount invested at 8%)
2500-x = 1000 dollars ( amount invested at 12%)
==============
Cheers,
Stan H.